Microorganism a) can biodegrade all organic compound b) extract energy and C atoms from biodegradable organics c) extract more energy if the organic is highly oxidized d) extract more energy if the organic is highly reduced e) are all harmful and disease-causing (pathogens) f) only, b and d are true The concentration of HOCL in water becomes exactly equal to that of OCL (its conjugate base) when: b- pHpKa c-pH = pKb e- c&d true f- a&b true For 99% inactivation of Fecal Coliform at a known pH and temperature, chlorine has: a- less C.t than Ozone b- more C.t than Ozone c- more lethality coefficient d- same C.t as Ozone 10. Salinity, as a water quality issue, can be measured by any of the following interdependent parameters except a) ionic strength (u) b) TDS c) TSS

environmental engineering
Answer #1

Answer 7: b) extract energy and C atoms from biodegradable organics

Answer 8: f) a and b true

Hint:pH=pKa+log\left [ OCl- \right ]/\left [ HOCL \right ]

POH=PKb+log\left [ HOCL \right ]/\left [ OCL- ^}\right ]

Where, concentration of HOCl and OCl- is equal so log 1=0.0 and thats why pH=PKa and POH=PKb